OpenStudy (anonymous):

Identify the 42nd term of an arithmetic sequence where a1 = -12 and a27 = 66

OpenStudy (lgbasallote):

let A_n = 66 so n= 27 66 = -12 + (27 -1)d 66 = -12 + (26d) 66 + 12 = 26d 78 = 26d d = 3 so to find A_42 use A1 = -12 n = 42 d = 3 can you do that now?

OpenStudy (anonymous):

Remember that\[a_{n} = a_{1} + (n-1)d\]where d is the common difference. Solve for d, and you will get 3. Then, use the same formula for n = 42:\[a_{42} = -12 + 41*3 = 111\]

OpenStudy (anonymous):

Identify the 27th term of an arithmetic sequence where a1 = 38 and a17 = -74.

OpenStudy (anonymous):

and thanks btw

OpenStudy (anonymous):

Use the same formula and thought process again and you will be gold :-)

OpenStudy (anonymous):

i got 96.5 :/ wat did i do wrong

OpenStudy (anonymous):

Maybe you solved for d wrongly. Recheck it. I got d = 7 for the common difference.

OpenStudy (anonymous):

A) -20.5 B) -151 C) -22.75 D) -144 those ar ethe possible choices

OpenStudy (anonymous):

Oops, I got d = -7. So, we get:\[a_{27} = a_1 + 26*(-7) = -144\]

OpenStudy (anonymous):

Identify the 35th term of an arithmetic sequence where a1 = -7 and a18 = 95 thanks